The peak power consumed by a resistive coil when connected to an AC source is 80 W. Find the energy consumed by the coil in 100 seconds which is many times larger than the time period of the source.

Grade:upto college level

1 Answers

Jitender Pal
9 years ago
Sol. P base 0 = 80 W (given) P base rms = P base 0/2 = 40 W Energy consumed = P * t = 40 * 100 = 4000 J = 40 KJ